# TWIN-GPT: Digital Twins for Clinical Trials via Large Language Model

### v1

@@ 5.3.1 Severe outcome prediction
We train an LSTM to predict the severe outcome. This model takes in the real clinical data and synthetic data generated by TWIN-GPT. We show this severe outcome prediction result in Fig 8. The LSTM achieves nearly the same AUROC scores when taking in the synthetic data generated by TWIN-GPT as taking in the real data, indicating that TWIN-GPT can fully understand and reproduce the latent relationship between records and severe outcome in real data.
@@ 5.3.2 Adverse event prediction
This task is to see how much causal relation the synthetic data have maintained compared to the raw clinical data. We train an MLP to predict the adverse event at the next time step, by taking in real and synthetic data generated by different models separately. Results are shown in Fig 8, where both TWIN-GPT and $k$-NN-based model have similar performance with the real data, indicating that our method captures the temporal causal relations accurately.
@@ 5.4.1 Presence Disclosure
If an attacker finds that the synthetic data was trained by TWIN-GPT from the patient n’s record $X_{n;1:T_{n}}$, we call it presence disclosure. We assume that there are total $m$of training data that had been known by the attackers, and set $m$ as 1%, 5%, 10%, and 20%. After calculating the sensitivity through Eq.7, we display the result in Fig 9(a). As the number of known samples rises, the sensitivity of TWIN-GPT has remained stable at around 20% (the attackers can recognize 20% of the real patients’ records from synthetic data they know). Correspondingly, the $k$-NN-based method reaches the maximum sensitivity of around 50%. Additionally, we observe that in terms of presence disclosure, TWIN-GPT’s sensitivity is slightly higher than that of TWIN-VAE but significantly lower than the $k$-NN algorithm. This could be because TWIN-GPT, as a large language model-based method, might possess a greater capability to capture and replicate the subtle differences of the training data. Such high fidelity in data replication may inadvertently lead to synthetic data retaining too many identifiable features of the original training data, thus increasing the risk of presence disclosure. However, on the TOP dataset, both the overall medical trial sensitivity for presence disclosure and the phase-by-phase medical trial sensitivity for presence disclosure are better than $k$-NN in Fig 10.
@@ 5.4.2 Attribute Disclosure
Here, we assume that the attackers know parts of the training set and have access to $x$of features of the records. We set $x$ as 1, 5, 10, and 20 in this experiment. When an attacker can infer additional attributes of an individual based on the features of a subset of the data he knows, the attribute disclosure occurs. We use Eq.8 to calculate mean sensitivity. The results are shown in Fig 9(b). We can see that the mean sensitivity of TWIN-GPT is generally less than the $k$-NN-based method and Twin method, meaning better performance. We can also observe that the maximum score of $k$-NN-based method and TWIN-GPT are respectively around 0.3 and 0.25. On the TOP dataset, as shown in Fig 11, the mean sensitivity of TWIN-GPT is significantly lower than that of the $k$-NN algorithm. Across different numbers of features, TWIN-GPT’s mean sensitivity is 0.2 percentage points lower than that of $k$-NN.
@@ 5.4.3 Nearest neighbor adversarial accuracy risk
NNAA is a measure of the degree to which a model overfits the original data, directly relating to privacy leakage risk. Here, we select 71 participants with 500 visit records as evaluation set $S_{E}$. Correspondingly, we also choose 500 clinical records from real data as training set $S_{T}$ and from synthetic data to form $S_{S}$, like the method required. TWIN-GPT achieves a score of 0.271. Generally, If the NNAA score is close to 0.5, we say that models overfit the original data. Therefore it suggests that TWIN-GPT is not excessively memorizing the original data and is instead learning more generalizable patterns.

### v2

@@ 5.3.1 Severe outcome prediction
We train an LSTM to predict the severe outcome. This model takes in the real clinical data and synthetic data generated by TWIN-GPT. We show this severe outcome prediction result in Fig. 8. The LSTM achieves nearly the same AUROC scores when taking in the synthetic data generated by TWIN-GPT as taking in the real data, indicating that TWIN-GPT can fully understand and reproduce the latent relationship between records and severe outcomes in real data.
@@ 5.3.2 Adverse event prediction
This task is to see how much causal relation the synthetic data have maintained compared to the raw clinical data. We train an MLP to predict the adverse event at the next step by taking in real and synthetic data generated by different models separately. Results are shown in Fig. 8, where both TWIN-GPT and $k$-NN-based model have similar performance with the real data, indicating that our method captures the temporal causal relations accurately.
@@ 5.4.1 Presence Disclosure
If an attacker finds that the synthetic data was trained by TWIN-GPT from the patient n’s record $X_{n;1:T_{n}}$, we call it presence disclosure. We assume that there are total $m$of training data that had been known by the attackers and set $m$ as 1%, 5%, 10%, and 20%. After calculating the sensitivity through Eq. 7, we display the result in Fig. 9(a). As the number of known samples rises, the sensitivity of TWIN-GPT has remained stable at around 20% (the attackers can recognize 20% of the real patients’ records from synthetic data they know). Correspondingly, the $k$-NN-based method reaches the maximum sensitivity of around 50%. Additionally, we observe that in terms of presence disclosure, TWIN-GPT’s sensitivity is slightly higher than that of TWIN-VAE but significantly lower than the $k$-NN algorithm. This could be because TWIN-GPT, as a large language model-based method, might possess a greater capability to capture and replicate the subtle differences of the training data. Such high fidelity in data replication may inadvertently lead to synthetic data retaining too many identifiable features of the original training data, thus increasing the risk of presence disclosure. However, on the TOP dataset, both the overall medical trial sensitivity for presence disclosure and the phase-by-phase medical trial sensitivity for presence disclosure are better than $k$-NN in Fig. 10.
@@ 5.4.2 Attribute Disclosure
Here, we assume that the attackers know parts of the training set and have access to $x$of features of the records. We set $x$ as 1, 5, 10, and 20 in this experiment. When an attacker can infer additional attributes of an individual based on the features of a subset of the data he knows, the attribute disclosure occurs. We use Eq. 8 to calculate mean sensitivity. The results are shown in Fig. 9(b). We can see that the mean sensitivity of TWIN-GPT is generally less than the $k$-NN-based method and Twin method, meaning better performance. We can also observe that the maximum score of $k$-NN-based method and TWIN-GPT are respectively around 0.3 and 0.25. On the TOP dataset, as shown in Fig. 11, the mean sensitivity of TWIN-GPT is significantly lower than that of the $k$-NN algorithm. Across different numbers of features, TWIN-GPT’s mean sensitivity is 0.2 percentage points lower than that of $k$-NN.
@@ 5.4.3 Nearest neighbor adversarial accuracy risk
NNAA is a measure of the degree to which a model overfits the original data, directly relating to privacy leakage risk. Here, we select 71 participants with 500 visit records as evaluation sets $S_{E}$. Correspondingly, we also choose 500 clinical records from real data as training sets $S_{T}$ and from synthetic data to form $S_{S}$, like the method required. TWIN-GPT achieves a score of 0.271. Generally, If the NNAA score is close to 0.5, we say that models overfit the original data. Therefore it suggests that TWIN-GPT is not excessively memorizing the original data and is instead learning more generalizable patterns.
@@ 5.6 Ablation Study
6 CONCLUSIONS
In this paper, leveraging ChatGPT as the base model, we developed a specific large language model called TWIN-GPT, for generating personalized digital twins tailored for virtual clinical trial simulation. The approach proves advantageous for efficient and accurate prediction of clinical trial outcomes, particularly when confronted with limited EHR training data. The TWIN-GPT demonstrates exceptional performance, especially in terms of fidelity, utility, and privacy. The evaluation further validates that the TWIN-GPT can pose low privacy risks concerning presence disclosure, attribute disclosure, and nearest neighbor adversarial accuracy risks, effectively addressing privacy concerns inherent in traditional physical clinical trials.